# RSCC: A Large-Scale Remote Sensing Change Caption Dataset for Disaster Events

### v2

@@ 1 Introduction
Temporal remote sensing imagery is indispensable for monitoring dynamic Earth processes, particularly disaster events that demand rapid response and analysis. Temporal remote sensing data has proven indispensable in supporting disaster relief planning and response (Rahnemoonfar et al., 2023, 2021; Gupta et al., 2019). However, the inherently complex spatiotemporal relationships embedded within this data pose significant challenges for effective analysis and interpretation.
Advancements in the modeling of multimodal data have enabled generalist Multimodal Large Language Models (MLLMs) (Anthropic AI, 2024a, b; OpenAI, 2024, 2023; Gemini Team, 2024b, a; Llama Team, 2024; Qwen Team, 2023; Wang et al., 2024d; Meta AI, ) that can perform a variety of natural image interpretation tasks specified flexibly through natural language. Specifically, MLLMs trained in a interleaved way have a deep visual-semantic understanding across images (Chen et al., 2024; Li et al., 2024a; Mistral AI, 2024; Wang et al., 2024c; Xue et al., 2024; Li et al., 2024b). These models achieve great success in multi-image reasoning (Meng et al., 2024; Wang et al., 2024b; Zhao et al., 2024) and video understanding (Liu et al., 2024c; Chandrasegaran et al., 2024; Cai et al., 2024; He et al., 2024; Fu et al., 2024), while their capabilities in temporal remote sensing image understanding remain underexplored.
Existing remote sensing image-text datasets often focus on single-snapshot imagery and lack the temporal details vital for understanding dynamic events, particularly in disaster-related scenarios, As shown in Table 1. Although there are multi-temporal datasets (e.g ., fMoW (Christie et al., 2018), SpaceNet 7 (Van Etten et al., 2021), S2Looking (Shen et al., 2021), QFabric (Verma et al., 2021) and SpaceNet 8 (Hänsch et al., 2022)), none of them provide rich textual descriptions of how scenes change over time. However, their potential for disaster-specific temporal analysis remains untapped due to the absence of high-quality bi-temporal datasets with detailed textual annotations. Existing remote sensing datasets either focus on generic land-use changes or provide short captions lacking disaster context. For instance, LEVIR-$C$ (Liu et al., 2022b) annotates urban development but omits disaster-specific details, while Dubai-CCD (Hoxha et al., 2022b) offers brief descriptions without capturing nuanced damage levels or infrastructure transformations.
@@
UCM-Captions (Qu et al., 2016) 2016 2,100 (1.0B) 10,500 (12)
RSICD (Lu et al., 2018) 2018 10,921 (0.5B) 54,605 (12)
fMoW (Christie et al., 2018) 2018 1M (437.0B) N/A
SpaceNet 7 (Van Etten et al., 2021) 2021 2,389 (2.6B) N/A
S2Looking (Shen et al., 2021) 2021 5,000 (5.0B) N/A
QFabric (Verma et al., 2021) 2021 2,520 (245.1B) N/A
SpaceNet 8 (Hänsch et al., 2022) 2022 2,576 (3.0B) N/A
LEVIR-$C$ (Liu et al., 2022b) 2022 20,154 (1.2B) 50,385 (40)
Dubai-CCD (Hoxha et al., 2022b) 2022 1,000 (<0.1B) 2,500 (35)
RSICap (Hu et al., 2023) 2023 2,585 (0.6B) 2,585 (60)
RS5M (Zhang et al., 2024) 2024 5M (-) 5M (49)
VRSBench (Li et al., ) 2024 29,614 (7.8B) 29,614 (52)
WHU-CDC (Shi et al., 2024) 2024 14,868 (1.9B) 37,170 (-)
XLRS-Bench (Wang et al., 2025) 2025 934 (67.5B) 934 (379)
@@ 2 Pipeline
To construct our RSCC dataset, we employed a multimodal reasoning model - Qwen QvQ-Max (Qwen Team, 2025b) - along with existing human label to generate high fidelity captions. QvQ-Max is the latest proprietary MLLM that is capable of visual reasoning which shows superior capabilities in zero-shot remote sensing image change caption (see Appendix A). Unlike traditional MLLMs that prioritize recognition-based outputs, QvQ-Max leverages a structured reasoning process to infer spatial-temporal relationships Bi et al. (2025). The QvQ-Max captioning process takes about \$5/k image pairs. The overall dataset construction pipeline is shown in Figure 2.
@@ 2.1 Data Source
In this study, we utilize xBD dataset (Gupta et al., 2019) along with EBD dataset (Wang et al., 2024f), which are all obtained from MAXAR OpenData Program. The images are cropped without overlapping to 512×512 from xBD’s original 1024×1024, while EBD retains its 512×512 resolution. The overall RSCC datasets consists of 62,351 bi-temporal pre- and post-disaster image pairs (xBD: 44,136; EBD: 18,215) spanning from 31 events covering disaster types ranging from earthquake, flooding (hurricane), tsunami, storm (hurricane, tornado), volcano eruption and wildfire. Full events list is shown in Appendix A.
@@ 2.2 Attribute Extraction
The xBD dataset contains human annotations of building bounding boxes with damage assessment labels. The damage evaluation is based on the Joint Damage Scale (Gupta et al., 2019), which was developed with contributions from organizations such as NASA and the California Air National Guard. This scale is designed to assess building damage from satellite imagery across various disaster scenarios, providing detailed descriptions for different level ranging from no damage to destroyed.
@@ 2.3 Prompt Construction
We carefully design the following instructions to prompt QvQ-Max Qwen Team (2025b) to create detailed bi-temporal image change captions. We convert building damage labels into in-context auxiliary information. Shtedritski et al. Shtedritski et al. (2023) found that by applying marking-based visual prompt engineering, it is possible to unlock effective behaviors in vision-language models like CLIP Radford et al. (2021), even without any training examples. This approach led to state-of-the-art results in zero-shot referring expression comprehension tasks. Inspired by this idea, we construct building damage masks as visual prompts for MLLMs.
@@ 2.5 Post-Correction and Human Verification
To ensure the reliability of captions generated by QvQ-Max, we implement a two-stage post-correction process. First, the Qwen2.5-Max (Qwen et al., ) systematically enforces metadata alignment by correcting discrepancies in disaster type (e.g., resolving mismatches between "hurricane" and metadata-specified "flooding") and damage descriptions (e.g., revising "minor damage" to "destroyed" based on building annotations). This automated stage achieves disaster type consistency with metadata from 98.7% to 100.0%. Second, a subset of RSCC captions (988 samples) is manually validated by three experts using a 0/1 binary rubric across four criteria: disaster type accuracy, damage detail completeness, factual consistency, and clarity. 92.3% of sampled captions passed validation, with 89.3% inter-annotator agreement (Cohen’s κ). Failed captions were reprocessed through the automated pipeline with refined rules, ensuring final dataset consistency. Full details of correction rules and evaluation protocols are provided in Appendix A.
@@ 3.2 RSCC for vision-language model training
In order to facilitate the vision-language model training, we divide the RSCC dataset into two splits, where a train split contains 61,327 image pairs from 31 distinct events across xBD as well as EBD and a test split contains 988 image pairs from 19 distinct events in xBD. We conduct full-parameter fine-tuning on Qwen2.5-VL 7B (Qwen Team, 2025a) using train set for 2 epoch with a batch_size=1batch\_size=1 on a single node equipped with 2 NVIDIA H800 GPUs. We initialize the learning rate at 1e-6 and 1e-5 for LLM backbone and vision encoder respectively and employ a cosine learning rate decay schedule for optimization. For image inputs, we maintain the native resolution of RSCC as 512×512 for maximum pixel inputs and minimum pixel inputs. The training procedure cost a total of 40 GPU clock hours.
@@ 4.1 Experiment Settings
For remote sensing change captioning, we benchmark moderate size open-sourced MLLMs (less than 13B parameters) that supports multi-image inputs, including LLaVA-NeXT-Interleave (Li et al., 2024b), xGen-MM [^2] https://hugggingfae.co/Salesforce/xgen-$m$-phi3-mini-instruct-interleave-r-v1.5 (BLIP-3) (Xue et al., 2024), LLaVA-OneVision (Li et al., 2024a), Qwen2-VL (Wang et al., 2024c),Pixtral (Mistral AI, 2024) Phi-4-Multimodal (Microsoft, 2025) Kimi-VL (Kimi Team, 2025) and InternVL 3 (Zhu et al., 2025). We also add two specialized remote sensing change captioning models namely TEOChat (Irvin et al., 2024) and CCExpert (Wang et al., 2024e).
@@
For model evaluation, we compare the text similarity with n-gram overlap metrics including ROUGE (Lin, 2004) and METEOR (Banerjee and Lavie, 2005). While the aforementioned measures are commonly reported in image captioning works, we find they are suboptimal to measure the semantic similarity across long texts. Therefore, we follow Kaggle LLM Prompt Recovery Competition  [^3] https://www.kaggle.com/c/llm-prompt-recovery and introduce Sentence T5-XXL Embedding (Ni et al., 2022) with Sharpened Cosine Similarity (Brandon, 2022) (ST5-SCS) to get a well-established similarity measure. We set $q=0$ and $p=3$ for sharpened cosine similarity.
@@ 4.2 Quantitative Results
BLIP-3 (3B) (Xue et al., 2024) 4.53 10.85 44.05 456
Kimi-VL (3B)(Kimi Team, 2025) 12.47 16.95 51.35 87
Phi-4-Multimodal (4B) (Microsoft, 2025) 4.09 1.45 34.55 7
Qwen2-VL (7B)(Wang et al., 2024c) 11.02 9.95 45.55 42
LLaVA-NeXT-Interleave (8B) (Li et al., 2024b) 12.51 13.29 46.99 57
LLaVA-OneVision (8B)(Li et al., 2024a) 8.40 10.97 46.15 221
InternVL 3 (8B) (Zhu et al., 2025) 12.76 15.77 51.84 64
Pixtral (12B) (Mistral AI, 2024) 12.34 15.94 49.36 70
CCExpert (7B) (Wang et al., 2024e) 7.61 4.32 40.81 12
TEOChat (7B)(Irvin et al., 2024) 7.86 5.77 52.64 15
@@
3. Repetition Issue BLIP-3 and LLaVA-OneVision are prone to generative repetitive outputs. It is assumed that these models fail in dealing with remote sensing images or following complex instructions. This degeneration problem may be alleviated by switching decoding methods (e.g., Contrastive Decoding Su et al. ) as well as adapting generation configurations Welleck et al. .
@@ 4.3 Human Preference Study
While the language metrics can be biased, we ask experts to vote the best caption from two anonymous model output given the bi-temporal image pairs along with human labeled building damage masks from xBD dataset Gupta et al. (2019). Results (Figure 4) reveal QvQ-Max (ground truth change captions) consistently outperformed all baselines, achieving win rates ranging from 80.7% (against InternVL3) to 99.0% (against CCExpert). While strong baselines like InternVL3 (19.3% wins) and mid-tier models (e.g., Pixtral [14.6%], Kimi-VL [12.8%]) showed moderate performance, our captions demonstrated superior accuracy in capturing fine-grained environmental changes critical for disaster response. Weak-performing multimodal baselines (LLaVA-Interleave [5.2%], Phi-4-$M$ [4.9%]) highlighted limitations in handling complex spatiotemporal reasoning, suggesting QvQ-Max’s quantization-aware training and dynamic context adaptation mechanisms enhance generalization. These findings validate QvQ-Max as a state-of-the-art solution for vision-language tasks in remote sensing.
@@ 4.4.2 Scaling Correction Decoding
To investigate the effectiveness of scaling correction decoding strategies (e.g., VCD Leng et al. (2024), DoLa Chuang et al. (2023) and DeCo Wang et al. (2024a)) in mitigating hallucinations during remote sensing change captioning, we evaluated their impact across varying model sizes for Qwen2.5-VL and InternVL3 (Figure 6). These strategies aim to align model outputs with input scale or context, reducing inconsistencies in multimodal reasoning.
@@ 5.1 Remote Sensing Change Captioning Models
The remote sensing image change captioning (RSICC) task aims to generate detailed and accurate natural language to describe geospatial feature changes for remote sensing images captured at different times (Chang and Ghamisi, 2023a; Hoxha et al., 2022a). Liu et al. (Liu et al., 2022a) introduced RSICCformer, a Transformer-based approach that incorporates multiple cross-encoding modules to leverage differential features, focusing attention on the changing regions within each image. Similarly, Chg2Cap Chang and Ghamisi (2023b) presents a Transformer-based caption generation model that translates the relationship between image embeddings and word embeddings into descriptive text. Recent researches tend to use pre-trained LLM for language generation. GeoLLaVA (Elgendy et al., 2024) use fine-tune pretrained temporal MLLMs (i.e. Video-LLaVA Lin et al. (2024) and LLaVA-NeXT-Video Li et al. (2024b)) for detecting temporal changes in geographical landscapes. CCExpert Wang et al. (2024e), which is developed based on LLaVA-OneVision (Li et al., 2024a), introduces a difference-focused integration component. This module is designed to identify multi-scale variations between bi-temporal images and merge them into the initial image context. TEOChat Irvin et al. (2024) applys a shared vision-encoder to agument the temporal understanding capability of LLaVA-1.5 Liu et al. (2024b). Despite the common architecture that intergrates a pre-trained LLM backbone and a vision encoder, Diffusion-RSCC (Yu et al., 2025) utilize a probabilistic diffusion model for RSICC that focus on pixel-level differences under long time span.
@@ 5.2 Remote Sensing Change Caption Dataset
Datasets that combine temporal and vision-language elements play a crucial role in training models to comprehend and merge temporal dynamics with linguistic information Liu et al. (2024a). Unlike the VQA dataset Irvin et al. (2024); Elgendy et al. (2024) that can be easily formatted through mask labels, the common practice of remote sensing change caption dataset is to further annotate existing change detection dataset with 5 sentences each image pairs, such as Dubai-CCD (Hoxha et al., 2022b), LEVIR-$C$ (Liu et al., 2023) and WHU-CDC (Shi et al., 2024). Given the generalization capability of commercial MLLMs, Wang et al.  Wang et al. (2024e) leveraged GPT-4o OpenAI (2024), using the explicit information provided by the change masks to generate detailed change descriptions.

### v4

@@ 1 Introduction
Temporal remote sensing imagery is indispensable for monitoring dynamic Earth processes, particularly disaster events that demand rapid response and analysis. Temporal remote sensing data has proven indispensable in supporting disaster relief planning and response [61, 60, 19]. However, the inherently complex spatiotemporal relationships embedded within this data pose significant challenges for effective analysis and interpretation.
Advancements in the modeling of multimodal data have enabled generalist Multimodal Large Language Models (MLLMs) [2, 3, 52, 51, 18, 17, 41, 56, 74, 44] that can perform a variety of natural image interpretation tasks specified flexibly through natural language. Specifically, MLLMs trained in a interleaved way have a deep visual-semantic understanding across images [11, 30, 46, 73, 78, 31]. These models achieve great success in multi-image reasoning [43, 71, 81] and video understanding [40, 8, 7, 21, 15], while their capabilities in temporal remote sensing image understanding remain underexplored.
Existing remote sensing image-text datasets often focus on single-snapshot imagery and lack the temporal details vital for understanding dynamic events, particularly in disaster-related scenarios, As shown in Table 1. Although there are multi-temporal datasets (e.g ., fMoW [12], SpaceNet 7 [68], S2Looking [63], QFabric [69] and SpaceNet 8 [20]), none of them provide rich textual descriptions of how scenes change over time. However, their potential for disaster-specific temporal analysis remains untapped due to the absence of high-quality bi-temporal datasets with detailed textual annotations. Existing remote sensing datasets either focus on generic land-use changes or provide short captions lacking disaster context. For instance, LEVIR-CC [36] annotates urban development but omits disaster-specific details, while Dubai-CCD [23] offers brief descriptions without capturing nuanced damage levels or infrastructure transformations.
@@
UCM-Captions [53] 2016 2,100 (1.0B) 10,500 (12)
RSICD [42] 2018 10,921 (0.5B) 54,605 (12)
fMoW [12] 2018 1M (437.0B) N/A
SpaceNet 7 [68] 2021 2,389 (2.6B) N/A
S2Looking [63] 2021 5,000 (5.0B) N/A
QFabric [69] 2021 2,520 (245.1B) N/A
SpaceNet 8 [20] 2022 2,576 (3.0B) N/A
LEVIR-CC [36] 2022 20,154 (1.2B) 50,385 (40)
Dubai-CCD [23] 2022 1,000 (<0.1B) 2,500 (35)
RSICap [24] 2023 2,585 (0.6B) 2,585 (60)
RS5M [80] 2024 5M (-) 5M (49)
VRSBench [32] 2024 29,614 (7.8B) 29,614 (52)
WHU-CDC [64] 2024 14,868 (1.9B) 37,170 (-)
XLRS-Bench [72] 2025 934 (67.5B) 934 (379)
@@ 2 Pipeline
To construct our RSCC dataset, we employed a multimodal reasoning model - Qwen QvQ-Max [58] - along with existing human label to generate high fidelity captions. QvQ-Max is the latest proprietary MLLM that is capable of visual reasoning which shows superior capabilities in zero-shot remote sensing image change caption (see Appendix A). Unlike traditional MLLMs that prioritize recognition-based outputs, QvQ-Max leverages a structured reasoning process to infer spatial-temporal relationships [5]. The QvQ-Max captioning process takes about \$5/k image pairs. The overall dataset construction pipeline is shown in Figure 2.
@@ 2.1 Data Source
In this study, we utilize xBD dataset [19] along with EBD dataset [76], which are all obtained from MAXAR OpenData Program. The images are cropped without overlapping to 512×512 from xBD’s original 1024×1024, while EBD retains its 512×512 resolution. The overall RSCC datasets consists of 62,351 bi-temporal pre- and post-disaster image pairs (xBD: 44,136; EBD: 18,215) spanning from 31 events covering disaster types ranging from earthquake, flooding (hurricane), tsunami, storm (hurricane, tornado), volcano eruption and wildfire. Full events list is shown in Appendix A.
@@ 2.2 Attribute Extraction
The xBD dataset contains human annotations of building bounding boxes with damage assessment labels. The damage evaluation is based on the Joint Damage Scale [19], which was developed with contributions from organizations such as NASA and the California Air National Guard. This scale is designed to assess building damage from satellite imagery across various disaster scenarios, providing detailed descriptions for different level ranging from no damage to destroyed.
@@ 2.3 Prompt Construction
We carefully design the following instructions to prompt QvQ-Max [58] to create detailed bi-temporal image change captions. We convert building damage labels into in-context auxiliary information. Shtedritski et al. [65] found that by applying marking-based visual prompt engineering, it is possible to unlock effective behaviors in vision-language models like CLIP [59], even without any training examples. This approach led to state-of-the-art results in zero-shot referring expression comprehension tasks. Inspired by this idea, we construct building damage masks as visual prompts for MLLMs.
@@ 2.5 Post-Correction and Human Verification
To ensure the reliability of captions generated by QvQ-Max, we implement a two-stage post-correction process. First, the Qwen2.5-Max [54] systematically enforces metadata alignment by correcting discrepancies in disaster type (e.g., resolving mismatches between "hurricane" and metadata-specified "flooding") and damage descriptions (e.g., revising "minor damage" to "destroyed" based on building annotations). This automated stage achieves disaster type consistency with metadata from 98.7% to 100.0%. Second, a subset of RSCC captions (988 samples) is manually validated by three experts using a 0/1 binary rubric across four criteria: disaster type accuracy, damage detail completeness, factual consistency, and clarity. 92.3% of sampled captions passed validation, with 89.3% inter-annotator agreement (Cohen’s ϰ). Failed captions were reprocessed through the automated pipeline with refined rules, ensuring final dataset consistency. Full details of correction rules and evaluation protocols are provided in Appendix A.
@@ 3.2 RSCC for vision-language model training
In order to facilitate the vision-language model training, we divide the RSCC dataset into two splits, where a train split contains 61,327 image pairs from 31 distinct events across xBD as well as EBD and a test split contains 988 image pairs from 19 distinct events in xBD. We conduct full-parameter fine-tuning on Qwen2.5-VL 7B [57] using train set for 2 epoch with a batch_size=1batch\_size=1 on a single node equipped with 2 NVIDIA H800 GPUs. We initialize the learning rate at 1e-6 and 1e-5 for LLM backbone and vision encoder respectively and employ a cosine learning rate decay schedule for optimization. For image inputs, we maintain the native resolution of RSCC as 512×512 for maximum pixel inputs and minimum pixel inputs. The training procedure cost a total of 40 GPU clock hours.
@@ 4.1 Experiment Settings
For remote sensing change captioning, we benchmark moderate size open-sourced MLLMs (less than 13B parameters) that supports multi-image inputs, including LLaVA-NeXT-Interleave [31], xGen-MM [^2] https://hugggingfae.co/Salesforce/xgen-mm-phi3-mini-instruct-interleave-r-v1.5 (BLIP-3) [78], LLaVA-OneVision [30], Qwen2-VL [73],Pixtral [46] Phi-4-Multimodal [45] Kimi-VL [27] and InternVL 3 [82]. We also add two specialized remote sensing change captioning models namely TEOChat [25] and CCExpert [75].
@@
For model evaluation, we compare the text similarity with n-gram overlap metrics including ROUGE [34] and METEOR [4]. While the aforementioned measures are commonly reported in image captioning works, we find they are suboptimal to measure the semantic similarity across long texts. Therefore, we follow Kaggle LLM Prompt Recovery Competition  [^3] https://www.kaggle.com/c/llm-prompt-recovery and introduce Sentence T5-XXL Embedding [48] with Sharpened Cosine Similarity [6] (ST5-SCS) to get a well-established similarity measure. We set $q=0$ and $p=3$ for sharpened cosine similarity.
@@ 4.2 Quantitative Results
BLIP-3 (3B) [78] 4.53 10.85 44.05 456
Kimi-VL (3B)[27] 12.47 16.95 51.35 87
Phi-4-Multimodal (4B) [45] 4.09 1.45 34.55 7
Qwen2-VL (7B)[73] 11.02 9.95 45.55 42
LLaVA-NeXT-Interleave (8B) [31] 12.51 13.29 46.99 57
LLaVA-OneVision (8B)[30] 8.40 10.97 46.15 221
InternVL 3 (8B) [82] 12.76 15.77 51.84 64
Pixtral (12B) [46] 12.34 15.94 49.36 70
CCExpert (7B) [75] 7.61 4.32 40.81 12
TEOChat (7B)[25] 7.86 5.77 52.64 15
@@
3. Repetition Issue BLIP-3 and LLaVA-OneVision are prone to generative repetitive outputs. It is assumed that these models fail in dealing with remote sensing images or following complex instructions. This degeneration problem may be alleviated by switching decoding methods (e.g., Contrastive Decoding [66]) as well as adapting generation configurations [77].
@@ 4.3 Human Preference Study
While the language metrics can be biased, we ask experts to vote the best caption from two anonymous model output given the bi-temporal image pairs along with human labeled building damage masks from xBD dataset [19]. Results (Figure 4) reveal QvQ-Max (ground truth change captions) consistently outperformed all baselines, achieving win rates ranging from 80.7% (against InternVL3) to 99.0% (against CCExpert). While strong baselines like InternVL3 (19.3% wins) and mid-tier models (e.g., Pixtral [14.6%], Kimi-VL [12.8%]) showed moderate performance, our captions demonstrated superior accuracy in capturing fine-grained environmental changes critical for disaster response. Weak-performing multimodal baselines (LLaVA-Interleave [5.2%], Phi-4-MM [4.9%]) highlighted limitations in handling complex spatiotemporal reasoning, suggesting QvQ-Max’s quantization-aware training and dynamic context adaptation mechanisms enhance generalization. These findings validate QvQ-Max as a state-of-the-art solution for vision-language tasks in remote sensing.
@@ 4.4.2 Scaling Correction Decoding
To investigate the effectiveness of scaling correction decoding strategies (e.g., VCD [29], DoLa [13] and DeCo [70]) in mitigating hallucinations during remote sensing change captioning, we evaluated their impact across varying model sizes for Qwen2.5-VL and InternVL3 (Figure 6). These strategies aim to align model outputs with input scale or context, reducing inconsistencies in multimodal reasoning.
@@ 5.1 Remote Sensing Change Captioning Models
The remote sensing image change captioning (RSICC) task aims to generate detailed and accurate natural language to describe geospatial feature changes for remote sensing images captured at different times [9, 22]. Liu et al. [35] introduced RSICCformer, a Transformer-based approach that incorporates multiple cross-encoding modules to leverage differential features, focusing attention on the changing regions within each image. Similarly, Chg2Cap [10] presents a Transformer-based caption generation model that translates the relationship between image embeddings and word embeddings into descriptive text. Recent researches tend to use pre-trained LLM for language generation. GeoLLaVA [14] use fine-tune pretrained temporal MLLMs (i.e. Video-LLaVA [33] and LLaVA-NeXT-Video [31]) for detecting temporal changes in geographical landscapes. CCExpert [75], which is developed based on LLaVA-OneVision [30], introduces a difference-focused integration component. This module is designed to identify multi-scale variations between bi-temporal images and merge them into the initial image context. TEOChat [25] applys a shared vision-encoder to agument the temporal understanding capability of LLaVA-1.5 [39]. Despite the common architecture that intergrates a pre-trained LLM backbone and a vision encoder, Diffusion-RSCC [79] utilize a probabilistic diffusion model for RSICC that focus on pixel-level differences under long time span.
@@ 5.2 Remote Sensing Change Caption Dataset
Datasets that combine temporal and vision-language elements play a crucial role in training models to comprehend and merge temporal dynamics with linguistic information [38]. Unlike the VQA dataset [25, 14] that can be easily formatted through mask labels, the common practice of remote sensing change caption dataset is to further annotate existing change detection dataset with 5 sentences each image pairs, such as Dubai-CCD [23], LEVIR-CC [37] and WHU-CDC [64]. Given the generalization capability of commercial MLLMs, Wang et al.  [75] leveraged GPT-4o [52], using the explicit information provided by the change masks to generate detailed change descriptions.